# Localization and interpolation of parabolic $L^{p}$ Neumann problems

### v1

@@ 3. Basic estimates
    $W_{i}\subset 4KQ$ for $K$ large enough as $W\cap 4Q\neq\emptyset$.
By the triangle inequality, we have (recall that $c_{v}=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W_{Q}}v\,dXdt=(v)_{W_{M}}$)
(4.24) $\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)-c_{v}|{\mathds{1}}_{4Q}(X,t)dXdt\leq\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)-(v)_{W}|{\mathds{1}}_{4Q}dXdt+\sum_{i=2}^{M}|(v)_{W_{i-1}}-(v)_{W_{i}}|,$
@@
This together with (4.24),(4.25), and property (1) above, we conclude that
$\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)-c_{v}|{\mathds{1}}_{4Q}(X,t)dXdt\lesssim\ln\left(\frac{r}{\ell(W)}\right)\left(\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})(y,s)+\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})(y,s)\right).$
@@
$\iint_{4Q\cap\Omega}|v-c_{v}|dXdt\lesssim r\int_{(y,s)\in\partial\Omega}\left(\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})(y,s)+\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})(y,s)\right)d\sigma(y,s),$
@@
Some additional observations are needed for estimating the 3 terms from (4.21). Notice that all these terms involve derivatives of $\varphi$, and so the regions of integrals are all contained in $2Q\cap\Omega\setminus\frac{3}{2}Q$. Since $G=G{\mathds{1}}_{Q}$, the function $v-c_{v}$ satisfies the homogeneous equation $L^{*}(v-c_{v})=0$ in $2Q\cap\Omega\setminus\frac{3}{2}Q$. We cover the region $2Q\cap\Omega\setminus\frac{3}{2}Q$ by a finite collection of parabolic cubes $Q_{i}^{\prime}$ of sidelength $\ell(Q_{i}^{\prime})\approx r$ such that
@@
    either $Q_{i}^{\prime}$ is centered at $\partial\Omega$ and $4Q_{i}^{\prime}\cap Q=\emptyset$,
@@
    or $4Q_{i}^{\prime}\subset 2Q\cap\Omega\setminus Q$.
Applying Lemma 3.3 to $v-c_{v}$ on $Q_{i}^{\prime}$ of type (i)(i)—recall that $v$ has 0 Neumann data on $\partial\Omega$—and applying the usual interior Moser estimate on $Q_{i}^{\prime}$ of type $(ii)$, we get that
(4.26) $\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega\setminus\frac{3}{2}Q}|v-c_{v}|^{2}dXdt\right)^{1/2}\leq\sup_{2Q\cap\Omega\setminus\frac{3}{2}Q}|v-c_{v}|\lesssim r^{-n-2}\iint_{4Q\cap\Omega\setminus Q}|v-c_{v}|dXdt.$
@@
(4.27) $\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega\setminus\frac{3}{2}Q}|v-c_{v}|^{2}dXdt\right)^{1/2}\lesssim r^{-n-1}\left(\|\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}\right)\\ \lesssim r^{-\frac{n+1}{p^{\prime}}}\left(\|\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})\|_{L^{p^{\prime}}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})\|_{L^{p^{\prime}}(\partial\Omega)}\right)\lesssim r^{-\frac{n+1}{p^{\prime}}},$
@@
Now we are ready to esimate the 3 terms from (4.21). For $I_{00}$, we use (4.20), (4.27), and the bound $|\partial_{t}\varphi|\lesssim r^{-2}$, to obtain that
$|I_{00}|\lesssim r^{-2}\sup_{2Q\cap\Omega}|u-c_{u}|\iint_{2Q\cap\Omega\setminus\frac{3}{2}Q}|v-c_{v}|dXdt\lesssim r^{-\frac{n+1}{p}}\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{16Q\cap\Omega}|\nabla u|dXdt.$
@@
$|I_{01}|\lesssim r^{-\frac{n+1}{p}}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega}|\nabla u|^{2}dXdt\right)^{1/2}.$
Finally, for $I_{02}$, we use (4.20), the bound on $|\nabla\varphi|$, Lemma 3.1 (as well as the decomposition of the region $2Q\cap\Omega\setminus\frac{3}{2}Q$ into parabolic cubes of types $(i)$ and $(ii)$), and (4.27) to get that
$|I_{02}|\lesssim r^{n+2}\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{16Q\cap\Omega}|\nabla u|dXdt\iint_{2Q\cap\Omega\setminus\frac{3}{2}Q}|\nabla v|dXdt\lesssim r^{-\frac{n+1}{p}}\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{16Q\cap\Omega}|\nabla u|dXdt.$
The estimate (4.17) follows as desired. ∎
@@
In the section, we present an application of the localization estimate that we have established in Theorem 1.4 and prove Theorem 1.6. Let
@@
We assume that, for some $p\in(1,\infty)$. both $\rm{(}\rm{N}_{p}\rm{)}$L andand $\rm{(}\rm{D}_{p^{\prime}}\rm{)}$$^{L^{*}}$ are solvable. We shall establish the end-point $p=1$ Hardy space bound of the form
@@
$\mbox{supp }g\subset Q_{r}(X,t)\cap\partial\Omega\quad\mbox{for some $(X,t)\in\partial\Omega$ and $r>0$ },\|g\|_{L^{\infty}}\leq r^{-n-2},\quad\int_{\partial\Omega}g=0.$
This bound implies solvability of the Neumann boundary value problem for $p=1$ with Hardy space data. Then by interpolation we get for all $1<q<p$ the solvability of the $\rm{(}\rm{N}_{q}\rm{)}$L, which proves Theorem 1.6. Here we interpolate the sublinear functional
@@
for some small $\eta>0$. To prove this we first establish that $\tilde{u}$ is bounded away from $Q_{1}(0,0)$. This follows from the assumption that the $L^{p}$ Neumann problem is solvable. In particular we have that $\|\tilde{N}(\nabla u)\|_{L^{p}(\partial\Omega)}<\infty$. For $(X,t)\in\mathbb{R}^{n}\times\mathbb{R}$, let $J_{r}(X,t)$ denote the backward in time parabolic cube, i.e.,
$J_{r}(X,t)=Q_{r}^{\prime}(X)\times(t-r^{2},t).$
@@
when $p>1$. Then by interior Poincare’s inequality for $\tilde{u}$ (Lemma 3.5)
@@
where $c_{\tilde{u}}$ denotes the average of $\tilde{u}$ in $J_{3/4}(X,t)$. Clearly, if we can take $c_{\tilde{u}}=0$ the claim would follow. We know however that this holds for t<1t<-1.We claim that simple geometric considerations imply that there exist $N=N(n)$ and a chain of regions $J_{1}(X_{1},t_{1}),J_{1}(X_{2},t_{2}),\dots,J_{1}(X_{k},t_{k})$ with $k\leq N$ such that
@@
When $X_{i}=X_{i+1}$ but $t_{i}\neq t_{i+1}$ the argument is a bit more involved and requires the use of PDE $\tilde{u}$ satisfies. This is done in full detail in [7] (see the calculation after (6.23) estimating the average on two parabolic cubes in term of $\nabla u$. We use the same idea in the proof of Lemma 3.5, see (3.7) and the subsequent calculation). Hence, indeed (5.4) holds for all points with $\delta(X,t)\leq O(1)$ and $t\leq O(1)$. Boundedness for the remaining points (i.e. those where either $\delta(X,t)$ is large or $t$ is large) follows by the maximum principle (which must hold even though our domain is unbounded in the $x_{n}$ variable, since the fact that $\tilde{u}$ is an energy solution implies a weak decay of $\tilde{u}$ when $|x_{n}|\to\infty$). Hence, $\tilde{u}$ is bounded on any complement of an enlargement of $Q_{1}(0,0)$.
@@
from which we deduce $\int_{\mathcal{\mathbb{R}}^{n}}\tilde{u}(X,10)dX=0$, as the extension $\tilde{u}$ is an even reflection of $u$. From this the claim (5.2) follows by applying the following lemma after substituting $t\mapsto t+10$ (c.f. Lemma 3.13 of [4]):
@@
(5.8) $\|\tilde{N}(\nabla u)\|_{L^{1}(Q_{16}(0,0)\cap\partial\Omega)}\leq|Q_{16}(0,0)\cap\partial\Omega|^{1/p^{\prime}}\left(\int_{Q_{16}(0,0)\cap\partial\Omega}\tilde{N}(\nabla u)^{p}d\sigma\right)^{1/p}\\ \lesssim|Q_{16}(0,0)\cap\partial\Omega|^{1/p^{\prime}}\left(\int_{\partial\Omega}g^{p}\right)^{1/p}\leq|Q_{16}(0,0)\cap\partial\Omega|^{1/p^{\prime}}|Q_{16}(0,0)\cap\partial\Omega|^{1/p}\\ =|Q_{16}(0,0)\cap\partial\Omega|\leq C,$

### v2

@@ 3. Basic estimates
    $W_{i}\subset 4KQ$ for $K$ large enough (as $W\cap 4Q\neq\emptyset$).
By the triangle inequality, we have (recall that $0=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W_{Q}}v\,dXdt=(v)_{W_{M}}$)
(4.25) $\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)|{\mathds{1}}_{4Q}(X,t)dXdt\leq\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)-(v)_{W}|{\mathds{1}}_{4Q}dXdt+\sum_{i=2}^{M}|(v)_{W_{i-1}}-(v)_{W_{i}}|,$
@@
This together with (4.25),(4.26), and property (1) above, allows us to conclude that
$\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{W}|v(X,t)|{\mathds{1}}_{4Q}(X,t)dXdt\lesssim\ln\left(\frac{r}{\ell(W)}\right)\left(\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})(y,s)+\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})(y,s)\right).$
@@
$\iint_{4Q\cap\Omega}|v|dXdt\lesssim r\int_{(y,s)\in\partial\Omega}\left(\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})(y,s)+\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})(y,s)\right)d\sigma(y,s),$
@@
Some additional observations are needed for estimating the 3 terms from (4.22). Notice that all these terms involve derivatives of $\varphi$, and so the regions of integrals are all contained in $J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}$. We cover the region $J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}$ by a finite collection of parabolic cubes $Q_{i}$ of sidelength $\ell(Q_{i})\approx r$ such that
@@
    either $Q_{i}$ is centered at $\partial\Omega$ and $4Q_{i}^{\prime}\cap J_{r}=\emptyset$,
@@
    or $4Q_{i}\subset\frac{5}{2}Q\cap\Omega\setminus J_{r}$.
Note that $v$ satisfies the homogeneous equation $L^{*}v=0$ in $\Omega\setminus J_{r}$ and has 0 Neumann data on $\partial\Omega$. Applying Lemma 3.3 to $v$ on $Q_{i}$ of type $(i)$ and applying the usual interior Moser estimate on $Q_{i}$ of type $(ii)$, we get that
(4.27) $\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}}|v|^{2}dXdt\right)^{1/2}\leq\sup_{J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}}|v|\lesssim r^{-n-2}\iint_{4Q\cap\Omega\setminus J_{r}}|v|dXdt.$
@@
(4.28) $\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}}|v|^{2}dXdt\right)^{1/2}\lesssim r^{-n-1}\left(\|\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})\|_{L^{1}(\partial\Omega)}\right)\\ \lesssim r^{-\frac{n+1}{p^{\prime}}}\left(\|\widetilde{N}(\delta|\nabla v|{\mathds{1}}_{4KQ})\|_{L^{p^{\prime}}(\partial\Omega)}+\|\widetilde{\mathcal{C}}(\delta G{\mathds{1}}_{4KQ})\|_{L^{p^{\prime}}(\partial\Omega)}\right)\lesssim r^{-\frac{n+1}{p^{\prime}}},$
@@
Now we are ready to estimate the 3 terms from (4.22). We choose $c_{u}=\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{8r}\cap\Omega}u\,dXdt$. Since $u-c_{u}$ is also a solution to $Lu=0$ in $Q_{0}\cap\Omega$ with zero Neumann data on $Q_{0}\cap\partial\Omega$, we have by Lemma 3.3 and Lemma 3.5 (recall $J_{Kr}\subset Q_{0}$) that
(4.29) $\sup_{J_{2r}\cap\Omega}|u-c_{u}|\leq C\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{8r}\cap\Omega}|u-c_{u}|dXdt\leq C\,r\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{16r}\cap\Omega}|\nabla u|dXdt.$
For $I_{00}$, we use (4.29), (4.28), and the bound $|\partial_{t}\varphi|\lesssim r^{-2}$ in $J_{2r}$, to obtain that
$|I_{00}|\lesssim r^{-2}\sup_{J_{2r}\cap\Omega}|u-c_{u}|\iint_{J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}}|v|dXdt\lesssim r^{-\frac{n+1}{p}}\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{16r}\cap\Omega}|\nabla u|dXdt.$
@@
$|I_{01}|\lesssim r^{-\frac{n+1}{p}}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega}|\nabla u|^{2}dXdt\right)^{1/2}.$
Finally, for $I_{02}$, we use (4.29), the bound on $|\nabla\varphi|$, Lemma 3.1 (as well as the decomposition of the region $J_{2r}\cap\Omega\setminus J_{\frac{2}{r}}$ into parabolic cubes of types $(i)$ and $(ii)$ as before), and (4.28) to get that
$|I_{02}|\lesssim r^{n+2}\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{16r}\cap\Omega}|\nabla u|dXdt\iint_{J_{2r}\cap\Omega\setminus J_{\frac{3}{2}r}}|\nabla v|dXdt\lesssim r^{-\frac{n+1}{p}}\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{16r}\cap\Omega}|\nabla u|dXdt.$
The estimate (4.17) follows as desired, and we can take $K=16$. ∎
@@
In this section, we present an application of the localization estimate that we have established in Theorem 1.4 and prove Theorem 1.6. Let
@@
We assume that, for some $p\in(1,\infty)$, both $\rm{(}\rm{N}_{p}\rm{)}$L andand $\rm{(}\rm{D}_{p^{\prime}}\rm{)}$$^{L^{*}}$ are solvable. We shall establish the end-point $p=1$ Hardy space bound of the form
@@
$\mbox{supp }g\subset Q_{r}(X,t)\cap\partial\Omega\quad\mbox{for some $(X,t)\in\partial\Omega$ and $r>0$ },\|g\|_{L^{\infty}}\leq r^{-n-1},\quad\int_{\partial\Omega}g=0.$
This bound implies solvability of the Neumann boundary value problem for $p=1$ with Hardy space data. Then by interpolation we get for all $1<q<p$ solvability of the $\rm{(}\rm{N}_{q}\rm{)}$L, which proves Theorem 1.6. Here we interpolate the sublinear functional
@@
for some small $\eta>0$. To prove this we first establish that $\tilde{u}$ is bounded away from $Q_{1}(0,0)$. This follows from the assumption that the $L^{p}$ Neumann problem is solvable. In particular we have that $\|\tilde{N}(\nabla u)\|_{L^{p}(\partial\Omega)}<\infty$.
@@
when $p>1$. Then by interior Poincaré’s inequality for $\tilde{u}$ (Lemma 3.5)
@@
where $c_{\tilde{u}}$ denotes the average of $\tilde{u}$ in $J_{3/4}(X,t)$. Clearly, if we can take $c_{\tilde{u}}=0$ the claim would follow. We know however that this holds for t<1t<-1. We claim that simple geometric considerations imply that there exist $N=N(n)$ and a chain of regions $J_{1}(X_{1},t_{1}),J_{1}(X_{2},t_{2})$$\dots,$ $J_{1}(X_{k},t_{k})$ with $k\leq N$ such that
@@
When $X_{i}=X_{i+1}$ but $t_{i}\neq t_{i+1}$ the argument is a bit more involved and requires the use of the PDE $\tilde{u}$ satisfies. This is done in full detail in [7] (see the calculation after (6.23) estimating the average on two parabolic cubes in term of $\nabla u$. We use the same idea in the proof of Lemma 3.5, see (3.7) and the subsequent calculation). Hence, indeed (5.4) holds for all points with $\delta(X,t)\leq O(1)$ and $t\leq O(1)$. Boundedness for the remaining points (i.e. those where either $\delta(X,t)$ is large or $t$ is large) follows by the maximum principle (which must hold even though our domain is unbounded in the $x_{n}$ variable, since the fact that $\tilde{u}$ is an energy solution implies a weak decay of $\tilde{u}$ when $|x_{n}|\to\infty$). Hence, $\tilde{u}$ is bounded on any complement of an enlargement of $Q_{1}(0,0)$.
@@
from which we deduce $\int_{\mathbb{R}^{n}}\tilde{u}(X,10)dX=0$, as the extension $\tilde{u}$ is an even reflection of $u$. From this the claim (5.2) follows by applying the following lemma after substituting $t\mapsto t+10$ (c.f. Lemma 3.13 of [4]):
@@
(5.8) $\|\tilde{N}(\nabla u)\|_{L^{1}(Q_{16}(0,0)\cap\partial\Omega)}\leq|Q_{16}(0,0)\cap\partial\Omega|^{1/p^{\prime}}\left(\int_{Q_{16}(0,0)\cap\partial\Omega}\tilde{N}(\nabla u)^{p}d\sigma\right)^{1/p}\\ \lesssim|Q_{16}(0,0)\cap\partial\Omega|^{1/p^{\prime}}\left(\int_{\partial\Omega}|g|^{p}\right)^{1/p}\leq|Q_{16}(0,0)\cap\partial\Omega|^{1/p^{\prime}}|Q_{16}(0,0)\cap\partial\Omega|^{1/p}\\ =|Q_{16}(0,0)\cap\partial\Omega|\leq C,$